# Sharp Analysis for KL-Regularized Contextual Bandits and RLHF

### v1

@@ 3.4 Discussion: Result for Local-Coverage
In comparison with the sample complexity $\Theta(\eta^{2}D^{2}+\eta/\epsilon)$ under global-coverage in Theorem 3.3, the order $\Theta(C_{\rho_{\mathrm{KL}}}\eta/\epsilon)$ depends on a weaker coverage $C_{\rho_{\mathrm{KL}}}$, but only has a multiplicative dependence on the coverage coefficient instead of additive dependence. whether the additive dependence can be achieved under the local-coverage condition is left as future work. Moreover, we compare this result with Theorem 4.2 in Song et al. (2024) and suppose that the in-sample-error $\epsilon_{\mathrm{reward}}$ of Song et al. (2024) is $O(1/n)$, their sample complexity is $\Theta(C_{\rho_{\mathrm{KL}}}^{2}/\epsilon^{2})$, which is looser than ours $\Theta(C_{\rho_{\mathrm{KL}}}\eta/\epsilon)$ when $\eta=o(C_{\rho_{\mathrm{KL}}}/\epsilon)$.
@@ 4 Reinforcement Learning from Preference Feedback
In this section, we consider the problem of aligning the language model with preference feedback. As discussed in Section 2.2, at each round, we can sample a pair of actions (responses) $a_{1},a_{2}$ and call a preference oracle to get the preference label $y\in\{0,1\}$, where $y=1$ means that the user prefers $a_{1}$ over $a_{2}$ (Definition 2.2). To learn the reward function, we introduce the following assumption for step 1 to ensure the existence of an MLE estimation oracle that can globally maximize the likelihood of the BT model over all possible reward functions.
@@
Following the previous analysis for RLHF (Xiong et al., 2024a), we assume the existence of a policy improvement oracle (Definition 2.5, corresponding to step 2) that can compute the optimal policy $\pi_{\widehat{\theta}}^{\eta}$ based on the reward function $\widehat{\theta}$.
@@ 4.1 Theoretical guarantees
For any $\epsilon\in(0,1),\eta>0$, and any algorithm $A$, there exists a KL-regularized preference learning problem as defined in Section 2.2 with $O(1)$ coverage coefficient and reward function class $\mathcal{R}$ such that $A$ requires at least $\Omega\big{(}\min(\frac{\eta\log N_{\mathcal{R}}(\epsilon)}{\epsilon},\frac{\log N_{\mathcal{R}}(\epsilon)}{\epsilon^{2}})\big{)}$ samples to achieve a suboptimality gap of $\epsilon$.
We defer Algorithm 2, a 2-stage mixed-policy sampling algorithm for RLHF with preference feedback, to Appendix A for conciseness because of its similarity to Algorithm 1.
Upper bound for global coverage
@@
We compare our two-stage mixed sampling method with hybrid frameworks. From the algorithmic perspective, a hybrid algorithm first learns from an offline dataset and then requires sufficient online iterations to ensure the performance (Xiong et al., 2024a). For example, for a finite action space with $A$ actions, the number of online iterations should be $\Theta(A)$. In contrast, our method only requires two iterations of sampling from mixed policy and interacting with the environment. Moreover, the results of hybrid literature depend on both the coverage coefficient and the structure complexity of the function class (like the dimension for a linear function class or eluder dimension (Russo and Van Roy, 2013)). Our result only needs the coverage condition of the reference policy. More importantly, we obtain a sharper bound on the sample complexity and derive the additive dependence on the coverage coefficient.
@@
Although the coefficient $e^{B}$ appearing in sample size $m,n$ can be exponentially large, this term is caused by the non-linearity of the link function for the preference model, and is common in RLHF literature (Zhu et al., 2023; Xiong et al., 2024a; Ye et al., 2024b; Song et al., 2024).
@@
5 Experimental Results
(a) Benefit of mixed-policy sampling
(b) Results under different KL-penalty
In this section, we conduct experiments with synthetic data to investigate the benefit of mixed-policy sampling and the effect of KL-regularization coefficient on the sample complexity of the problem. We plot the experimental results for RL from preference feedback in Figure 1 and defer the results for KL-regularized contextual bandits in Appendix B. All the trials are repeated for 10 times and plotted with the standard variation.
We consider the case where context distribution $d_{0}$ is a projected Gaussian distribution over the unit sphere and $\mathcal{A}$ is a discrete set with $|\mathcal{A}|=5$. We construct the reward functions as $R(\phi,x,a)=\langle x,\phi(a)\rangle$, parameterized by a mapping $\phi$ from $\mathcal{A}$ to $\mathbb{R}^{10}$. To generate $\phi_{*}$, we sample $\phi_{*}(a)$ independently for each $a\in\mathcal{A}$ according to another projected gaussian distribution over the sphere with radius equal to 5. In Figure 1(a), we compare the suboptimality gaps of mixed-policy sampling with $m=n$ to those of offline sampling using $\pi_{0}$ under the same sample sizes. The result indicates that the usage of mixed-policy sampling reduces the suboptimality gap by a large margin. In Figure 1(b), it is shown that the sample complexity is remarkably affected by the KL-regularization term, corroborating our sharp analysis for regularized RLHF.
6 Conclusion
We have presented a comprehensive theoretical analysis of the role of reverse-KL regularization in decision-making models including contextual bandits and reinforcement learning from preference feedback, highlighting its significance in terms of sample complexity. Our results provide new insights into the power of regularization extending beyond its traditional role of mitigating errors from the current critic (or reward) model.
Additionally, we examined the role of data coverage in both contextual bandits and RLHF. Our analysis shows that with sufficient coverage from the reference policy, a mixed sampling strategy can achieve a sample complexity that exhibits only an additive dependence on the coverage coefficient without the need for explicit exploration or additional structural assumptions.
@@ Acknowledgements
Appendix A Algorithm for Preference Feedback
Algorithm 2 2-Stage mixed-policy sampling for preference feedback
@@ Appendix A Algorithm for Preference Feedback
Appendix B Additional Experimental Results
@@
We show our experimental results for KL-regularized contextual bandits in the following figure 2, which also corroborate our theory in Section 3.
@@ Appendix C Proofs from Section 3
Proof of Theorem 3.1.
Consider a simple case when $|\mathcal{X}|=M$ and $|\mathcal{A}|=2$. We suppose that the context $x$ is drawn uniformly from $\mathcal{X}$ at the beginning of each round. Let $\Theta$ be the set consisting of mappings from $\mathcal{X}$ to $\mathcal{A}=\{0,1\}$. For each $\theta\in\Theta$, we have $R(\theta,x,a)=\begin{cases}1/2+c&\text{if }a=\theta(x),\\ 1/2&\text{if }a\neq\theta(x),\end{cases}$ where $c>0$ is a constant, and $\theta(x)$ is the optimal action under context $x$ when the model is $\theta$.
@@
$\leq 2\cdot\Bigl{(}\frac{1}{2}\log\frac{1}{1+2c}+\frac{1}{2}\log\frac{1}{1-2c}\Bigr{)}=-\log(1-4c^{2}).$
@@
$|\mathbb{P}_{\theta_{1}}(A)-\mathbb{P}_{\theta_{2}}(A)|\leq\sqrt{-\frac{1}{2}\log(1-4c^{2})\mathbb{E}_{\theta_{1}}[N(x_{0})]}\leq\sqrt{2c^{2}\mathbb{E}_{\theta_{1}}[N(x_{0})]}=\sqrt{2c^{2}T/M},$
@@
Set $A$ to be the event that $\pi_{out}(\theta_{1}(x_{0})|x_{0})>1/2$. Then we have
$\mathbb{P}_{\theta_{1}}(\pi_{out}(\theta_{1}(x_{0})|x_{0})\leq 1/2)+\mathbb{P}_{\theta_{2}}(\pi_{out}(\theta_{2}(x_{0})|x_{0})\leq 1/2)\geq 1-|\mathbb{P}_{\theta_{1}}(A)-\mathbb{P}_{\theta_{2}}(A)|\geq 1-\sqrt{2c^{2}T/M}.$
@@
$\mathbb{E}_{\theta\sim Unif(\Theta)}\mathbb{P}_{\theta}(\pi_{out}(\theta(x_{0}))\leq 1/2)\geq\frac{1}{2}-\sqrt{c^{2}T/2M}$
@@
$\mathbb{E}_{\pi_{\theta_{*}}^{\eta}}\bigg{[}R(\theta_{*},x,a)-\frac{1}{\eta}\ln\frac{\pi_{\theta_{*}}^{\eta}(a|x)}{\pi_{0}(a|x)}\bigg{]}-\mathbb{E}_{\pi_{out}}\bigg{[}R(\theta_{*},x,a)-\frac{1}{\eta}\ln\frac{\pi_{out}(a|x)}{\pi_{0}(a|x)}\bigg{]}$
$=\frac{1}{\eta}\mathbb{E}_{\pi_{\theta_{*}}^{\eta}}\bigg{[}\ln\frac{\pi_{0}(a|x)\cdot\exp\bigl{(}\eta R(\theta_{*},x,a)\bigr{)}}{\pi_{\theta_{*}}^{\eta}(a|x)}\bigg{]}-\frac{1}{\eta}\mathbb{E}_{\pi_{out}}\bigg{[}\ln\frac{\pi_{0}(a|x)\cdot\exp\bigl{(}\eta R(\theta_{*},x,a)\bigr{)}}{\pi_{out}(a|x)}\bigg{]}$
$=\frac{1}{\eta}\mathbb{E}_{\pi_{out}}\Bigl{[}\ln\frac{\pi_{out}(a|x)}{\pi^{*}(a|x)}\Bigr{]},$
@@
$\mathbb{E}_{\theta\sim Unif(\Theta)}\mathbb{E}_{\pi_{out}}\Bigl{[}\ln\frac{\pi_{out}(a|x)}{\pi^{*}(a|x)}\Bigr{]}$
$=\mathbb{E}_{\theta\sim Unif(\Theta)}\frac{1}{M}\sum_{x\in\mathcal{X}}\mathbb{E}_{a\sim\pi_{out}(\cdot|x)}\Bigl{[}\ln\frac{\pi_{out}(a|x)}{\pi^{*}(a|x)}\Bigr{]}$
$\geq\mathbb{E}_{\theta\sim Unif(\Theta)}\frac{1}{M}\sum_{x\in\mathcal{X}}\mathbb{P}_{\theta}(\pi_{out}(\theta(x))\leq 1/2)\cdot\biggl{[}\frac{1}{2}\ln\frac{1+\exp(-\eta c)}{2}+\frac{1}{2}\ln\frac{1+\exp(\eta c)}{2}\biggr{]}$
$\geq\Bigl{(}\frac{1}{2}-\sqrt{c^{2}T/2M}\Bigr{)}\Bigl{[}\frac{1}{2}\ln\frac{1+\exp(-\eta c)}{2}+\frac{1}{2}\ln\frac{1+\exp(\eta c)}{2}\Bigr{]}$ (C.1)
Note that
$\frac{\text{d}}{\text{d}u}\Bigl{[}\frac{1}{2}\ln\frac{1+e^{-u}}{2}+\frac{1}{2}\ln\frac{1+e^{u}}{2}\Bigr{]}\Big{|}_{u=0}$ $=\frac{1}{2}\Bigl{[}\frac{1}{1+\exp(-u)}-\frac{1}{1+\exp(u)}\Bigr{]}\Big{|}_{u=0}=0,$
$\frac{\text{d}^{2}}{\text{d}u^{2}}\Bigl{[}\frac{1}{2}\ln\frac{1+e^{-u}}{2}+\frac{1}{2}\ln\frac{1+e^{u}}{2}\Bigr{]}$ $=\frac{\exp(u)}{[1+\exp(u)]^{2}}.$
@@
$\mathbb{E}_{\theta\sim Unif(\Theta)}\mathbb{E}_{\pi_{out}}\Bigl{[}\ln\frac{\pi_{out}(a|x)}{\pi^{*}(a|x)}\Bigr{]}\geq\frac{1}{2}\cdot\Bigl{(}\frac{1}{2}-\sqrt{c^{2}T/2M}\Bigr{)}\eta^{2}c^{2}\cdot\frac{1}{3+\exp(\eta c)},$
@@
$\frac{1}{2}\cdot\Bigl{(}\frac{1}{2}-\sqrt{c^{2}T/2M}\Bigr{)}\eta^{2}c^{2}\cdot\frac{1}{3+\exp(\eta c)}\leq\eta\epsilon,$
@@
$\frac{1}{2}\ln\frac{1+\exp(-\eta c)}{2}+\frac{1}{2}\ln\frac{1+\exp(\eta c)}{2}$ $=\frac{1}{2}\ln\frac{2+\exp(\eta c)+\exp(-\eta c)}{4}$
$\geq\frac{1}{2}\cdot\frac{1}{2}\Bigl{(}\ln\frac{\exp(\eta c)+\exp(-\eta c)}{2}\Bigr{)}$
$\geq\frac{1}{4}(\eta c-\ln 2),$ (C.2)
@@
$\epsilon\geq\frac{1}{\eta}\mathbb{E}_{\theta\sim Unif(\Theta)}\mathbb{E}_{\pi_{out}}\Bigl{[}\ln\frac{\pi_{out}(a|x)}{\pi^{*}(a|x)}\Bigr{]}\geq\frac{1}{4}\cdot\Bigl{(}\frac{1}{2}-\sqrt{c^{2}T/2M}\Bigr{)}(\eta c-\ln 2)\cdot\frac{1}{\eta}.$
@@
Proof of Theorem 3.3.

### v2

@@ 3.5 Result under Local-Coverage Condition
In comparison with the sample complexity $\Theta(\eta^{2}D^{2}+\eta/\epsilon)$ under data coverage in Theorem 3.8, the order $\Theta(C_{\rho_{\mathrm{KL}}}\eta/\epsilon)$ depends on a local coverage coefficient $C_{\rho_{\mathrm{KL}}}$, but has a multiplicative dependence on the coverage coefficient instead of additive dependence. whether the additive dependence can be achieved under the local-coverage condition is left as future work. Moreover, we compare this result with Theorem 4.2 in Song et al. (2024) and suppose that the in-sample-error $\epsilon_{\mathrm{reward}}$ of Song et al. (2024) is $O(1/n)$, their sample complexity is $\Theta(C_{\rho_{\mathrm{KL}}}^{2}/\epsilon^{2})$, which is looser than ours $\Theta(C_{\rho_{\mathrm{KL}}}\eta/\epsilon)$ when $\eta=o(C_{\rho_{\mathrm{KL}}}/\epsilon)$.
@@ 4 Reinforcement Learning from Preference Feedback
In this section, we consider the problem of aligning the language model with preference feedback.
4.1 Problem Setup
In each round, we can sample a pair of actions (responses) $a_{1},a_{2}$ and query a preference oracle to get the preference label $y\in\{0,1\}$, where $y=1$ means that the user prefers $a_{1}$ over $a_{2}$. Specifically, when receiving a prompt $x\in\mathcal{X}$, and two actions (responses) $a^{1},a^{2}\in\mathcal{A}$ from some LLM policy $\pi(\cdot|x)$, a preference oracle will give feedback $y$ defined as follows:
Definition 4.1 (Preference Oracle).
Definition 4.2 (Bradley-Terry Model).
where $\sigma(u)=(1+e^{-u})^{-1}$ is the sigmoid function.
To learn the reward function, we introduce the following assumption to ensure the existence of an MLE estimation oracle that can globally maximize the likelihood of the BT model over all possible reward functions.
@@ 4.1 Problem Setup
Following the previous analysis for RLHF (Xiong et al., 2024a), we assume the existence of a policy improvement oracle (Definition 3.4) that can compute the optimal policy $\pi_{\widehat{\theta}}^{\eta}$ based on the reward function $\widehat{\theta}$.
@@
In RLHF setting, we cannot directly observe or estimate absolute reward values. Consequently, the most intuitive estimation approach is to focus on relative rewards: for any context $x$ and actions $a^{1},a^{2}$, our estimated difference $f(s,a^{1})-f(s,a^{2})$ should closely approximate the true reward difference $r(s,a^{1})-r(s,a^{2})$. Therefore, we extend the data coverage condition to the RLHF setting as follows.
Definition 4.5 (Data Coverage).
Given a reference policy $\pi_{0}$, $D^{2}$ is the minimum positive real number satisfying $\forall\ {(x,a)\in\mathcal{X}\times\mathcal{A},\ \pi(a|x)>0}$, we have for any pair of $\theta,\theta^{\prime}\in\Theta$, there exists $b:\mathcal{X}\to[-B,B]$ such that
$\frac{|R(\theta^{\prime},x,a)-R(\theta,x,a)-b(x)|^{2}}{\mathbb{E}_{x^{\prime}\sim d_{0}}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi_{0}(\cdot|x^{\prime})}[R(\theta^{\prime},x^{\prime},a^{\prime})-R(\theta,x^{\prime},a^{\prime})]}\leq D^{2}.$
We also use the linear model to describe this condition, where the reward function $R(\theta,x,a)=\theta^{\top}\phi(x,a)$ for $\theta\in\mathbb{R}^{d}$. Let the covariance matrix
$\widetilde{\Sigma}=\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a\sim\pi_{0}(\cdot|x)}(\phi(x,a)-\mathbb{E}_{a^{\prime}\sim\pi_{0}(\cdot|x)}\phi(x,a^{\prime}))(\phi(x,a)-\mathbb{E}_{a^{\prime}\sim\pi_{0}(\cdot|x)}\phi(x,a^{\prime}))^{\top}.$
the coverage condition means that for any pair of $\theta,\theta^{\prime}$, there exist $b(x)=\theta^{\top}\mathbb{E}_{a^{\prime}\sim\pi_{0}(\cdot|x)}\phi(x,a^{\prime})$ such that
$\sup_{\theta,\theta^{\prime}\in\Theta}\frac{|(\theta^{\prime}-\theta)^{\top}\phi(x,a)-b(x)|^{2}}{(\theta^{\prime}-\theta)^{\top}\Sigma(\theta^{\prime}-\theta)}\leq\|\phi(x,a)-\mathbb{E}_{a^{\prime}\sim\pi_{0}(\cdot|x)}\phi(x,a^{\prime})\|_{\Sigma^{-1}}^{2},$
where the inequality uses the Cauchy-Schwarz inequality. This term can be bounded by $D^{2}=O(d)$ through G-optimal design (Zhang, 2023; Lattimore and Szepesvári, 2020).
@@ 4.2 Theoretical Guarantees
For any $\epsilon\in(0,1/256),\eta>4$, and any algorithm $A$, there exists a KL-regularized preference learning problem with $O(N_{\mathcal{R}}(\epsilon))$ coverage coefficient and reward function class $\mathcal{R}$ such that $A$ requires at least $\Omega\big{(}\min(\frac{\eta\log N_{\mathcal{R}}(\epsilon)}{\epsilon},\frac{\log N_{\mathcal{R}}(\epsilon)}{\epsilon^{2}})\big{)}$ samples to achieve a suboptimality gap of $\epsilon$.
We present a two-stage mixed-policy sampling algorithm for RLHF, which can be seen as an extention of Algorithm 1. Due to space limit, we defer it to Algorithm 2 in Appendix A.
Upper bound
@@
We compare our two-stage mixed sampling method with hybrid RL. From the algorithmic perspective, a hybrid RL algorithm first learns from an offline dataset and then requires sufficient online iterations to ensure the performance (Xiong et al., 2024a). For example, for a finite action space with $A$ actions, the number of online iterations should be $\Theta(A)$. In contrast, our method only requires two iterations of sampling from mixed policy and interacting with the environment. Moreover, the results of hybrid literature depend on both the coverage coefficient and the structure complexity of the function class (like the dimension for a linear function class or eluder dimension (Russo and Van Roy, 2013)). Our result only needs the coverage condition of the reference policy. More importantly, we obtain a sharper bound on the sample complexity and derive the additive dependence on the coverage coefficient.
@@
Although the coefficient $e^{B}$ appearing in sample size $m,n$ can be exponentially large, this term is caused by the non-linearity of the link function for the preference model, and is common in RLHF literature (Zhu et al., 2023; Xiong et al., 2024a; Ye et al., 2024a; Song et al., 2024).
@@
5 Conclusions and Future Work
We have presented a comprehensive theoretical analysis of the role of reverse-KL regularization in decision-making models including contextual bandits and reinforcement learning from preference feedback, highlighting its significance in terms of sample complexity. Our results provide new insights into the power of regularization extending beyond its traditional role of mitigating errors from the current critic (or reward) model. Additionally, we examined the role of data coverage in both contextual bandits and RLHF. Our analysis shows that with sufficient coverage from the reference policy, a mixed sampling strategy can achieve a sample complexity that exhibits only an additive dependence on the coverage coefficient without the need for explicit exploration or additional structural assumptions.
For future directions, it is interesting to study if the sharp bound exists for the KL-regularized Markov Decision Process (MDP) problem. Additionally, it is also worthwhile to explore how to achieve a sample complexity bound with only additive dependence on the global and local coverage conditions in Definitions 3.10 and 3.11.
@@ Acknowledgements
Appendix A KL-Regularized Algorithm for RLHF
In this section, we present a two-stage mixed-policy sampling algorithm for RLHF in Algorithm 2, which can be seen as an extention of Algorithm 1. There are two stages in the algorithm.
Algorithm 2 Two-stage Mixed-Policy Sampling from Preference Feedback (TMPS-PF)
@@ Appendix A KL-Regularized Algorithm for RLHF
Appendix B Experimental Results
@@
(a) Benefit of mixed-policy sampling
(b) Results under different KL-penalty
In this section, we conduct experiments with synthetic data to investigate the benefit of mixed-policy sampling and the effect of KL-regularization coefficient on the sample complexity of the problem. We plot the experimental results for RL from preference feedback in Figure 2 and the results for KL-regularized contextual bandits in Figure 1. All the trials are repeated for 10 times and plotted with the standard variation.
We consider the case where context distribution $d_{0}$ is a projected Gaussian distribution over the unit sphere and $\mathcal{A}$ is a discrete set with $|\mathcal{A}|=5$. We construct the reward functions as $R(\phi,x,a)=\langle x,\phi(a)\rangle$, parameterized by a mapping $\phi$ from $\mathcal{A}$ to $\mathbb{R}^{10}$, and set the reference policy $\pi_{0}$ to be the uniform random policy. To generate $\phi_{*}$, we sample $\phi_{*}(a)$ independently for each $a\in\mathcal{A}$ according to another projected gaussian distribution over the sphere with radius equal to 5. In Figure 2(a), we compare the suboptimality gaps of mixed-policy sampling with $m=n$ to those of offline sampling using $\pi_{0}$ under the same sample sizes. The result indicates that the usage of mixed-policy sampling reduces the suboptimality gap by a large margin. In Figure 2(b), it is shown that the sample complexity is remarkably affected by the KL-regularization term, corroborating our sharp analysis for regularized RLHF.
@@ Appendix C Proofs from Section 3
Proof.
Consider a simple case when $|\mathcal{X}|=M$ and $|\mathcal{A}|=2$. We suppose that the context $x$ is drawn uniformly from $\mathcal{X}$ at the beginning of each round. Let $\Theta$ be the set consisting of mappings from $\mathcal{X}$ to $\mathcal{A}=\{0,1\}$. For each $\theta\in\Theta$, we have $R(\theta,x,a)=\begin{cases}1/2+c&\text{if }a=\theta(x),\\ 1/2&\text{if }a\neq\theta(x),\end{cases}$ where $c\in(0,1/4)$ is a constant, and $\theta(x)$ is the optimal action under context $x$ when the model is $\theta$.
@@
$=2c\cdot\log\frac{1+2c}{1-2c}\leq 16c^{2}$
where the inequality follows from the fact that $\log(1+x)\leq x$ for $x\geq 0$ and $c\in(0,1/4)$.
@@
$|\mathbb{P}_{\theta_{1}}(A)-\mathbb{P}_{\theta_{2}}(A)|\leq\sqrt{8c^{2}\mathbb{E}_{\theta_{1}}[N(x_{0})]}=\sqrt{8c^{2}T/M},$
@@
Set $A$ to be the event that $\pi_{\text{out}}(\theta_{1}(x_{0})|x_{0})>1/2$. Then we have
$\mathbb{P}_{\theta_{1}}(\pi_{\text{out}}(\theta_{1}(x_{0})|x_{0})\leq 1/2)+\mathbb{P}_{\theta_{2}}(\pi_{\text{out}}(\theta_{2}(x_{0})|x_{0})\leq 1/2)\geq 1-|\mathbb{P}_{\theta_{1}}(A)-\mathbb{P}_{\theta_{2}}(A)|\geq 1-\sqrt{8c^{2}T/M}.$
@@
$\mathbb{E}_{\theta\sim\text{Unif}(\Theta)}\mathbb{P}_{\theta}(\pi_{\text{out}}(\theta(x_{0}))\leq 1/2)\geq\frac{1}{2}-\sqrt{2c^{2}T/M}$
@@
$\mathbb{E}_{\pi_{\theta_{*}}^{\eta}}\bigg{[}R(\theta_{*},x,a)-\frac{1}{\eta}\log\frac{\pi_{\theta_{*}}^{\eta}(a|x)}{\pi_{0}(a|x)}\bigg{]}-\mathbb{E}_{\pi_{\text{out}}}\bigg{[}R(\theta_{*},x,a)-\frac{1}{\eta}\log\frac{\pi_{\text{out}}(a|x)}{\pi_{0}(a|x)}\bigg{]}$
$=\frac{1}{\eta}\mathbb{E}_{\pi_{\theta_{*}}^{\eta}}\bigg{[}\log\frac{\pi_{0}(a|x)\cdot\exp\bigl{(}\eta R(\theta_{*},x,a)\bigr{)}}{\pi_{\theta_{*}}^{\eta}(a|x)}\bigg{]}-\frac{1}{\eta}\mathbb{E}_{\pi_{\text{out}}}\bigg{[}\log\frac{\pi_{0}(a|x)\cdot\exp\bigl{(}\eta R(\theta_{*},x,a)\bigr{)}}{\pi_{\text{out}}(a|x)}\bigg{]}$
$=\frac{1}{\eta}\mathbb{E}_{\pi_{\text{out}}}\Bigl{[}\log\frac{\pi_{\text{out}}(a|x)}{\pi^{*}(a|x)}\Bigr{]},$
@@
$\mathbb{E}_{\theta\sim\text{Unif}(\Theta)}\mathbb{E}_{\pi_{\text{out}}}\Bigl{[}\log\frac{\pi_{\text{out}}(a|x)}{\pi^{*}(a|x)}\Bigr{]}$
$=\mathbb{E}_{\theta\sim\text{Unif}(\Theta)}\frac{1}{M}\sum_{x\in\mathcal{X}}\mathbb{E}_{a\sim\pi_{\text{out}}(\cdot|x)}\Bigl{[}\log\frac{\pi_{\text{out}}(a|x)}{\pi^{*}(a|x)}\Bigr{]}$
$\geq\mathbb{E}_{\theta\sim\text{Unif}(\Theta)}\frac{1}{M}\sum_{x\in\mathcal{X}}\mathbb{P}_{\theta}(\pi_{\text{out}}(\theta(x))\leq 1/2)\cdot\biggl{[}\frac{1}{2}\log\frac{1+\exp(-\eta c)}{2}+\frac{1}{2}\log\frac{1+\exp(\eta c)}{2}\biggr{]}$
$\geq\Bigl{(}\frac{1}{2}-\sqrt{2c^{2}T/M}\Bigr{)}\Bigl{[}\frac{1}{2}\log\frac{1+\exp(-\eta c)}{2}+\frac{1}{2}\log\frac{1+\exp(\eta c)}{2}\Bigr{]},$ (C.1)
where the first inequality follows from the fact that $\mathrm{KL}(\pi_{\text{out}}(\cdot|x)\|\pi^{*}(\cdot|x))\geq\mathrm{KL}(\pi_{\text{unif}}(\cdot|x)\|\pi^{*}(\cdot|x))$ if $\pi_{\text{out}}(\theta(x))\leq 1/2$. Here $\pi_{\text{unif}}$ is the uniform distribution over $\mathcal{A}$. Note that
$\frac{\text{d}}{\text{d}u}\Bigl{[}\frac{1}{2}\log\frac{1+e^{-u}}{2}+\frac{1}{2}\log\frac{1+e^{u}}{2}\Bigr{]}\Big{|}_{u=0}$ $=\frac{1}{2}\Bigl{[}\frac{1}{1+\exp(-u)}-\frac{1}{1+\exp(u)}\Bigr{]}\Big{|}_{u=0}=0,$
$\frac{\text{d}^{2}}{\text{d}u^{2}}\Bigl{[}\frac{1}{2}\log\frac{1+e^{-u}}{2}+\frac{1}{2}\log\frac{1+e^{u}}{2}\Bigr{]}$ $=\frac{\exp(u)}{[1+\exp(u)]^{2}}.$
@@
$\mathbb{E}_{\theta\sim\text{Unif}(\Theta)}\mathbb{E}_{\pi_{\text{out}}}\Bigl{[}\log\frac{\pi_{\text{out}}(a|x)}{\pi^{*}(a|x)}\Bigr{]}\geq\frac{1}{2}\cdot\Bigl{(}\frac{1}{2}-\sqrt{2c^{2}T/M}\Bigr{)}\eta^{2}c^{2}\cdot\frac{1}{3+\exp(\eta c)},$
@@
$\frac{1}{2}\cdot\Bigl{(}\frac{1}{2}-\sqrt{2c^{2}T/M}\Bigr{)}\eta^{2}c^{2}\cdot\frac{1}{3+\exp(\eta c)}\leq\eta\epsilon,$
indicating that $T\geq\frac{\eta M}{2048\epsilon}=\Omega(\frac{\eta M}{\epsilon})$.
@@
$\frac{1}{2}\log\frac{1+\exp(-\eta c)}{2}+\frac{1}{2}\log\frac{1+\exp(\eta c)}{2}$ $=\frac{1}{2}\log\frac{2+\exp(\eta c)+\exp(-\eta c)}{4}$
$\geq\frac{1}{2}\cdot\frac{1}{2}\Bigl{(}\log\frac{\exp(\eta c)+\exp(-\eta c)}{2}\Bigr{)}$
$\geq\frac{1}{4}(\eta c-\log 2),$ (C.2)
@@
$\epsilon\geq\frac{1}{\eta}\mathbb{E}_{\theta\sim Unif(\Theta)}\mathbb{E}_{\pi_{\text{out}}}\Bigl{[}\log\frac{\pi_{\text{out}}(a|x)}{\pi^{*}(a|x)}\Bigr{]}\geq\frac{1}{4}\cdot\Bigl{(}\frac{1}{2}-\sqrt{c^{2}T/2M}\Bigr{)}(\eta c-\log 2)\cdot\frac{1}{\eta}.$
@@
Note that for the given $\epsilon$, $N_{\mathcal{R}}(\epsilon)=M$, and since $d_{0}$ is the uniform distribution over $\mathcal{X}$, we have $D^{2}=O(M)$, which completes the proof.
@@
Lemma C.5 (Restatement of Lemma 3.9).
For any estimator $\widehat{\theta}\in\Theta$, and the policy $\pi_{\widehat{\theta}}^{\eta}$ satisfying Definiton 3.4, we have
where $\Delta(x,a)=R(\widehat{\theta},x,a)-R(\theta_{*},x,a)$, $f(\cdot,\cdot)=\gamma R(\widehat{\theta},\cdot,\cdot)+(1-\gamma)R(\theta_{*},\cdot,\cdot)~{}(\gamma\in(0,1))$ the inequality uses the fact that second term on the right-hand side of the equality is $(\sum_{a\in\mathcal{A}}\pi_{f}^{\eta}(a|x)\Delta(x,a))^{2}\geq 0$.
Proof of Lemma 3.9.